An Eskimo returning from a successful friction trip pulls a sled loaded with salmon. The total mass of the sled and salmon is 50.0kg and the Eskimo exert a force of 120N on the sled by pulling on the rope.
a) How much work does he do on the sled if the rope is horizontal to the ground and he pulls the sled 5.00m ?
b)How much work does he do on the sled if tita equal 30.0° and he pulls the sled same distance
1
Expert's answer
2019-09-19T09:47:52-0400

By the definition of the work done we have:

a) "W = Fdcos \theta = 120 N \cdot 5.0 m \cdot cos0^{\circ} = 600 J."

b) "W = Fdcos \theta = 120 N \cdot 5.0 m \cdot cos30^{\circ} = 520 J."

Answer:

a) "W = 600 J."

b) "W = 520 J."